# Delensing for Precision Cosmology: Optimizing Future CMB B-mode Surveys to Constrain $r$

### v2

@@ 4.6.2 A generalized modeling result
We propose a more general model that additionally accounts for imperfections in the construction of the lensing B-mode template, parameterized by $A_{L}^{\prime}$. Given the complexity of the LT construction process, we expect the bias in the LT to be encapsulated by $A_{L}^{\prime}$, with $A_{L}^{\prime}=A_{L}=1$ representing a perfect construction in standard $\Lambda$CDM. Here, $A_{L}$ describes the lensed B-modes in the observation. Thus, deviations of $A_{L}$ from 1 ($A_{L}\neq 1$) will capture discrepancies between the observation and the theoretical model, while deviations of $A_{L}^{\prime}$ from 1 ($A_{L}^{\prime}\neq 1$) will reflect biases introduced by the pipeline.
@@
$C_{\ell}^{\text{SAT},\nu\nu^{\prime}}&=rC_{\ell}^{\text{tens}}|_{r=1}+A_{L}C_{\ell}^{\text{lens}}+N_{\ell}^{BB,\nu\nu^{\prime}}\\ &\quad+C_{\ell}^{\nu\nu^{\prime}}|_{\text{FG (7 params.)}},\end{aligned}$ (51)
@@
Generalized Model Baseline Model Parameter Fiducial value Before adding LT After adding LT (Internal) After adding LT (Combined tracer) After adding LT (Internal) After adding LT (Combined tracer) $r(\times 10^{3})$ 0 $-0.055\pm 1.800$ $-0.042\pm 1.048$ $-0.042\pm 0.686$ $-0.038\pm 0.958$ $-0.030\pm 0.649$ $A_{L}$ 1 $1.000\pm 0.049$ $1.000\pm 0.042$ $1.000\pm 0.040$ $1.000\pm 0.042$ $1.000\pm 0.037$ $A_{L}^{{}^{\prime}}$ 1 N/A $1.001\pm 0.074$ $1.000\pm 0.045$ N/A N/A $A_{d}(\mu K^{2})$ 14.300 $14.216\pm 0.640$ $14.215\pm 0.636$ $14.222\pm 0.634$ $14.215\pm 0.637$ $14.222\pm 0.633$ $\alpha_{d}$0.650 $-0.642\pm 0.073$ $-0.642\pm 0.072$ $-0.642\pm 0.072$ $-0.642\pm 0.072$ $-0.643\pm 0.072$ $\beta_{d}$ 1.48 $1.480\pm 0.010$ $1.480\pm 0.009$ $1.480\pm 0.007$ $1.480\pm 0.009$ $1.480\pm 0.007$ $A_{s}(\mu K^{2})$ 2.400 $2.378\pm 0.175$ $2.378\pm 0.170$ $2.378\pm 0.167$ $2.378\pm 0.170$ $2.378\pm 0.167$ $\alpha_{s}$0.800 $-0.775\pm 0.157$ $-0.777\pm 0.152$ $-0.779\pm 0.148$ $-0.778\pm 0.152$ $-0.778\pm 0.149$ $\beta_{s}$3.100 $-3.102\pm 0.035$ $-3.102\pm 0.031$ $-3.101\pm 0.028$ $-3.101\pm 0.031$ $-3.101\pm 0.028$ $\epsilon_{ds}(\times 10^{2})$ 0 $0.003\pm 3.811$ $0.015\pm 3.734$ $0.025\pm 3.677$ $0.003\pm 3.729$ $0.012\pm 3.684$
@@ A.1 Basic theory of NILC
$\hat{X}_{\text{NILC}}(\hat{\mathbf{n}})=\sum_{\ell m}\sum_{jk}\beta_{jk}^{\mathrm{NILC}}\sqrt{\frac{4\pi}{\lambda_{j}}}h_{\ell}^{j}Y_{\ell m}(\xi_{j})Y_{\ell m}(\hat{\mathbf{n}}).$ (61)
@@ B.1 Varying Cosmology
In this work, component maps are simulated based on various models. The simulation of CMB and lensing (and LSS tracers) is based on the $\Lambda$CDM model with parameters given by the Planck 2018 best-fit aghanim2020planck . The simulation of instrumental effects is described by the white noise level and the Gaussian beam. The simulation of foregrounds is based on models provided by PySM3 zonca2021python ; thorne2017python and the parameterization described in Section 3.3. In realistic observations, the difference between the simulated and observed data may introduce additional systematic effects. The impact of these effects on the delensing procedure should be carefully assessed to avoid introducing large bias on $r$ constraint.
@@
Planck 2018 ”TT,TE,EE+lowE+lensing+BAO” Varying Parameter Cases Params. 68% interval Exact Scale Underest. Large Scale Underest. Small Scale Overest. Small Scale Overest. Large Scale $\Omega_{b}h^{2}$ $0.02242\pm 0.00014$ $0.02242$ $0.02256$ $0.02256$ $0.02228$ $0.02228$ $\Omega_{c}h^{2}$ $0.11933\pm 0.00091$ $0.11933$ $0.11842$ $0.11842$ $0.12024$ $0.12024$ $100\theta_{\text{MC}}$ $1.04101\pm 0.00029$ $1.04101$ $1.04130$ $1.04130$ $1.04072$ $1.04072$ $\tau$ $0.0561\pm 0.0071$ $0.0561$ $0.0632$ $0.0632$ $0.0490$ $0.0490$ $ln(10^{1}0A_{s})$ $3.047\pm 0.014$ $3.047$ $3.033$ $3.033$ $3.061$ $3.061$ $n_{s}$ $0.9665\pm 0.0038$ $0.9665$ $0.9703$ $0.9627$ $0.9703$ $0.9627$
We consider varying six fundamental parameters for the base-$\Lambda$CDM model: $\Omega_{b}h^{2}$, $\Omega_{c}h^{2}$, $100\theta_{\text{MC}}$, $\tau$, $\ln(10^{10}A_{s})$, $n_{s}$}, within the 68% confidence interval as provided in the ”TT,TE,EE+lowE+lensing+BAO” case of Table 2 in aghanim2020planck , where $\Omega_{b}h^{2}$ is the physical baryon density, $\Omega_{c}h^{2}$ is the physical cold dark matter density, $\theta_{\text{MC}}$ is the angular acoustic scale (approximate), $\tau$ is the optical depth to reionization, $A_{s}$ is the amplitude of the primordial scalar power spectrum and $n_{s}$ represents the scalar spectral index. For each parameter, we assign four values between the 68% interval lower bound and upper bound. The six-dimensional parameter space consists of $4^{6}$ mesh points. We compute the lensing potential power spectrum and the lensed CMB power spectra at each of these points using CAMB lewis2011camb , resulting in $4^{6}$ curves, illustrated as the shaded region in Fig. 19.
@@ B.2 Effect on the Pipeline
Figure 19: The power spectra of the lensing potential, lensed TT, lensed EE, and lensed BB fields are computed using the parameters from a six-dimensional parameter space with $4^{6}$ mesh points. We show the ratio of the power spectra between the values at the mesh points and those at the best-fit parameters. The thick solid black lines delineate the upper and lower bounds of the ratio regions, and we have arbitrarily chosen five sets of parameter combinations for the simulations. For the lensed EE power, and the power below $\ell\sim 20$ is significantly affected by variations in the optical depth $\tau$ due to reionization.
@@
Underestimate Case Overestimate Case Parameter Input value Exact Case Small Scale Large Scale Small Scale Large Scale $r(\times 10^{3})$ 0 $-0.036\pm 0.954$ $0.073\pm 0.954$ $0.081\pm 0.957$ $-0.145\pm 0.955$ $-0.133\pm 0.953$ $A_{L}$ 1 $1.000\pm 0.043$ $0.999\pm 0.046$ $0.999\pm 0.046$ $1.001\pm 0.041$ $1.001\pm 0.041$ $A_{d}(\mu K^{2})$ 14.300 $14.214\pm 0.636$ $14.218\pm 0.638$ $14.219\pm 0.637$ $14.219\pm 0.638$ $14.212\pm 0.635$ $\alpha_{d}$0.650 $-0.642\pm 0.072$ $-0.641\pm 0.072$ $-0.643\pm 0.072$ $-0.643\pm 0.072$ $-0.643\pm 0.072$ $\beta_{d}$ 1.48 $1.480\pm 0.008$ $1.480\pm 0.008$ $1.480\pm 0.008$ $1.480\pm 0.008$ $1.480\pm 0.008$ $A_{s}(\mu K^{2})$ 2.400 $2.378\pm 0.170$ $2.380\pm 0.170$ $2.380\pm 0.170$ $2.377\pm 0.170$ $2.378\pm 0.170$ $\alpha_{s}$0.800 $-0.778\pm 0.152$ $-0.774\pm 0.152$ $-0.775\pm 0.152$ $-0.787\pm 0.152$ $-0.780\pm 0.152$ $\beta_{s}$3.100 $-3.102\pm 0.031$ $-3.102\pm 0.031$ $-3.102\pm 0.031$ $-3.102\pm 0.031$ $-3.102\pm 0.031$ $\epsilon_{ds}(\times 10^{2})$ 0 $-0.022\pm 3.736$ $-0.048\pm 3.719$ $-0.017\pm 3.740$ $-0.014\pm 3.736$ $-0.025\pm 3.730$
@@ C.1 Internal Lensing Reconstruction
    $\bar{X}(\hat{\mathbf{n}})=[\mathcal{B}^{\dagger}\text{Cov}^{-1}X^{\text{dat}}](\hat{\mathbf{n}}),$ (63)
@@
    A Gaussian likelihood is sufficient to describe the $\Lambda$CDM CMB lensing potential $\phi$, according to carron2017maximum , where an optimal Bayesian estimate is obtained by maximizing the log-posterior, implemented through iteration starting from a quadratic estimator (QE). In this paper, we focus on the QE, as we do not expect significant improvement from the Bayesian estimate for our experimental configuration. The spin-1 real-space (unnormalized) lensing displacement QE is given by
    ${{}_{1}}\hat{d}(\hat{\mathbf{n}})=-\sum_{s=0,\pm 2}{{}_{-s}}\bar{X}(\hat{\mathbf{n}})\big{[}\sharp_{s}X^{\text{WF}}\big{]}(\hat{\mathbf{n}}).$ (64)
@@
[…]
@@ Appendix D Some algorithms of delensing effect
$X^{de}(\hat{\mathbf{n}})$ $\approx\tilde{X}(\hat{\mathbf{n}}+\mathbf{d^{inv}_{1}}+\mathbf{d^{inv}_{2}})$ (73)
@@
$\nabla[(Q_{1}+Q_{2})\pm i(U_{1}+U_{2})](\hat{\mathbf{n}})\nabla[\phi_{1}(\hat{\mathbf{n}})+\phi_{2}(\hat{\mathbf{n}})]$ (76)
$=\nabla[Q_{1}\pm iU_{1}](\hat{\mathbf{n}})\nabla[\phi_{1}(\hat{\mathbf{n}})]+\nabla[Q_{2}\pm iU_{2}](\hat{\mathbf{n}})\nabla[\phi_{2}(\hat{\mathbf{n}})]$
$\quad+\nabla[Q_{2}\pm iU_{2}](\hat{\mathbf{n}})\nabla[\phi_{1}(\hat{\mathbf{n}})]+\nabla[Q_{1}\pm iU_{1}](\hat{\mathbf{n}})\nabla[\phi_{2}(\hat{\mathbf{n}})].$
@@ Appendix F A toy test of the generalized model
$\hat{C}_{\ell}^{\text{SAT},\nu\nu^{\prime}}&=1.02*C_{\ell}^{\text{lens}}+N_{\ell}^{BB,\nu\nu^{\prime}}+C_{\ell}^{\nu\nu^{\prime}}|_{\text{FG}},\end{aligned}$ (80)
@@
Generalized Model Baseline Model Parameter Fiducial value Before adding LT After adding LT (Combined tracer) After adding LT (Combined tracer) $r(\times 10^{3})$ 0 $-0.068\pm 1.808$ $-0.041\pm 0.691$ $0.198\pm 0.648$ $A_{L}$ 1.020 $1.020\pm 0.049$ $1.020\pm 0.040$ $1.005\pm 0.037$ $A_{L}^{{}^{\prime}}$ 0.980 N/A $0.980\pm 0.045$ N/A $A_{d}(\mu K^{2})$ 14.300 $14.213\pm 0.640$ $14.217\pm 0.633$ $14.206\pm 0.636$ $\alpha_{d}$0.650 $-0.642\pm 0.072$ $-0.642\pm 0.072$ $-0.643\pm 0.072$ $\beta_{d}$ 1.48 $1.480\pm 0.010$ $1.480\pm 0.007$ $1.480\pm 0.007$ $A_{s}(\mu K^{2})$ 2.400 $2.378\pm 0.175$ $2.378\pm 0.167$ $2.385\pm 0.167$ $\alpha_{s}$0.800 $-0.777\pm 0.157$ $-0.778\pm 0.148$ $-0.771\pm 0.148$ $\beta_{s}$3.100 $-3.102\pm 0.035$ $-3.102\pm 0.028$ $-3.101\pm 0.028$ $\epsilon_{ds}(\times 10^{2})$ 0 $0.015\pm 3.823$ $0.011\pm 3.694$ $0.010\pm 3.684$

### v3

@@ 4.6.2 A generalized modeling result
[…]
@@
$\begin{aligned} C_{\ell}^{\text{SAT},\nu\nu^{\prime}}&=rC_{\ell}^{\text{tens}}|_{r=1}+A_{L}C_{\ell}^{\text{lens}}+N_{\ell}^{BB,\nu\nu^{\prime}}\\ &\quad+C_{\ell}^{\nu\nu^{\prime}}|_{\text{FG (7 params.)}},\end{aligned}$ (51)
@@
Generalized Model Baseline Model Parameter Fiducial value Before adding LT After adding LT (Internal) After adding LT (Combined tracer) After adding LT (Internal) After adding LT (Combined tracer) $r(\times 10^{3})$ 0 $055\pm 1.800 0.0$$42\pm 1.048 0.0$$42\pm 0.686 0.0$$38\pm 0.958 0.0$$030\pm 0.649$ $A_{L}$ 1 $1.000\pm 0.049$ $1.000\pm 0.042$ $1.000\pm 0.040$ $1.000\pm 0.042$ $1.000\pm 0.037$ $A_{L}^{{}^{\prime}}$ 1 N/A $1.001\pm 0.074$ $1.000\pm 0.045$ N/A N/A $A_{d}(\mu K^{2})$ 14.300 $14.216\pm 0.640$ $14.215\pm 0.636$ $14.222\pm 0.634$ $14.215\pm 0.637$ $14.222\pm 0.633$ $\alpha_{d}$ -0.650 $0.642\pm 0.073$ $0.642\pm 0.072$ $0.642\pm 0.072$ $0.642\pm 0.072$ $0.643\pm 0.072$ $\beta_{d}$ 1.48 $1.480\pm 0.010$ $1.480\pm 0.009$ $1.480\pm 0.007$ $1.480\pm 0.009$ $1.480\pm 0.007$ $A_{s}(\mu K^{2})$ 2.400 $2.378\pm 0.175$ $2.378\pm 0.170$ $2.378\pm 0.167$ $2.378\pm 0.170$ $2.378\pm 0.167$ $\alpha_{s}$ -0.800 $0.775\pm 0.157$ $0.777\pm 0.152$ $0.779\pm 0.148$ $0.778\pm 0.152$ $0.778\pm 0.149$ $\beta_{s}$ -3.100 $3.102\pm 0.035$ $3.102\pm 0.031$ $3.101\pm 0.028$ $3.101\pm 0.031$ $3.101\pm 0.028$ $\epsilon_{ds}(\times 10^{2})$ 0 $003\pm 3.811 0.0$$15\pm 3.734 0.0$$25\pm 3.677 0.0$$03\pm 3.729 0.0$$012\pm 3.684$
@@ A.1 Basic theory of NILC
[…]
@@ B.1 Varying Cosmology
[…]
@@
Planck 2018 ”TT,TE,EE+lowE+lensing+BAO” Varying Parameter Cases Params. 68% interval Exact Scale Underest. Large Scale Underest. Small Scale Overest. Small Scale Overest. Large Scale $\Omega_{b}h^{2}$ $\pm 0.00014 0.022420.02242$ $0.02256$ $0.02256$ $0.02228$ $0.02228$ $\Omega_{c}h^{2}$ $\pm 0.00091 0.119330.11933$ $0.11842$ $0.11842$ $0.12024$ $0.12024$ $100\theta_{\text{MC}}$ $\pm 0.00029 1.041011.04101$ $1.04130$ $1.04130$ $1.04072$ $1.04072$ $\tau$ $\pm 0.0071 0.05610.0561$ $0.0632$ $0.0632$ $0.0490$ $0.0490$ $ln(10^{1}0A_{s})$ $\pm 0.014 3.0473.047$ $3.033$ $3.033$ $3.061$ $3.061$ $n_{s}$ $\pm 0.0038 0.96650.9665$ $0.9703$ $0.9627$ $0.9703$ $0.9627$
[…]
@@ B.2 Effect on the Pipeline
Figure 19: The power spectra of the lensing potential, lensed $T$, lensed $E$, and lensed BB fields are computed using the parameters from a six-dimensional parameter space with $4^{6}$ mesh points. We show the ratio of the power spectra between the values at the mesh points and those at the best-fit parameters. The thick solid black lines delineate the upper and lower bounds of the ratio regions, and we have arbitrarily chosen five sets of parameter combinations for the simulations. For the lensed $E$ power, and the power below $\ell\sim 20$ is significantly affected by variations in the optical depth $\tau$ due to reionization.
@@
Underestimate Case Overestimate Case Parameter Input value Exact Case Small Scale Large Scale Small Scale Large Scale $r(\times 10^{3})$ 0 $036\pm 0.954 0.0$$73\pm 0.954 0.0$$081\pm 0.957$ $0.145\pm 0.955$ $0.133\pm 0.953$ $A_{L}$ 1 $1.000\pm 0.043$ $0.999\pm 0.046$ $0.999\pm 0.046$ $1\pm 0.041 1.001±0.0$$1\pm 0.041$ $A_{d}(\mu K^{2})$ 14.300 $14.214\pm 0.636$ $14.218\pm 0.638$ $14.219\pm 0.637$ $14.219\pm 0.638$ $14.212\pm 0.635$ $\alpha_{d}$ -0.650 $0.642\pm 0.072$ $0.641\pm 0.072$ $0.643\pm 0.072$ $0.643\pm 0.072$ $0.643\pm 0.072$ $\beta_{d}$ 1.48 $1.480\pm 0.008$ $1.480\pm 0.008$ $1.480\pm 0.008$ $1.480\pm 0.008$ $1.480\pm 0.008$ $A_{s}(\mu K^{2})$ 2.400 $2.378\pm 0.170$ $2.380\pm 0.170$ $2.380\pm 0.170$ $2.377\pm 0.170$ $2.378\pm 0.170$ $\alpha_{s}$ -0.800 $0.778\pm 0.152$ $0.774\pm 0.152$ $0.775\pm 0.152$ $0.787\pm 0.152$ $0.780\pm 0.152$ $\beta_{s}$ -3.100 $3.102\pm 0.031$ $3.102\pm 0.031$ $3.102\pm 0.031$ $3.102\pm 0.031$ $3.102\pm 0.031$ $\epsilon_{ds}(\times 10^{2})$ 0 $022\pm 3.736 0.0$$48\pm 3.719 0.0$$17\pm 3.740 0.0$$14\pm 3.736 0.0$$025\pm 3.730$
@@ C.1 Internal Lensing Reconstruction
[…]
@@
[…]
@@
    where, $\Delta C^{\hat{\phi}_{1}\hat{\phi}_{2}}_{L}|_{\text{RDN0}}$ comes from disconnected signal from Gaussian fluctuation, and it exists even if the CMB maps do not have lensing in them, and it is of the zeroth order of $C^{\phi\phi}_{L}$. $\Delta C^{\hat{\phi}_{1}\hat{\phi}_{2}}_{L}|_{\text{N1}}$ comes from the connected part of 4-point function, which arises from the effects from other $\ell$ modes when we estimate a certain $\phi_{L}$, and it is of the first order of $C^{\phi\phi}_{L}$. The $N^{(1)}$ bias shows up as excess power $\left(\sim 10\%\right$. of $C_{L}^{\phi\phi}$ at $\left.L\sim 1000\right)$ on small scales.
@@ Appendix D Some algorithms of delensing effect
[…]
@@
[…]
@@ Appendix F A toy test of the generalized model
$\begin{aligned} \hat{C}_{\ell}^{\text{SAT},\nu\nu^{\prime}}&=1.02*C_{\ell}^{\text{lens}}+N_{\ell}^{BB,\nu\nu^{\prime}}+C_{\ell}^{\nu\nu^{\prime}}|_{\text{FG}},\end{aligned}$ (80)
@@
Generalized Model Baseline Model Parameter Fiducial value Before adding LT After adding LT (Combined tracer) After adding LT (Combined tracer) $r(\times 10^{3})$ 0 $068\pm 1.808 0.0$$041\pm 0.691$ $0.198\pm 0.648$ $A_{L}$ 1.020 $1.020\pm 0.049$ $1.020\pm 0.040$ $1.005\pm 0.037$ $A_{L}^{{}^{\prime}}$ 0.980 N/A $0.980\pm 0.045$ N/A $A_{d}(\mu K^{2})$ 14.300 $14.213\pm 0.640$ $14.217\pm 0.633$ $14.206\pm 0.636$ $\alpha_{d}$ -0.650 $0.642\pm 0.072$ $0.642\pm 0.072$ $0.643\pm 0.072$ $\beta_{d}$ 1.48 $1.480\pm 0.010$ $1.480\pm 0.007$ $1.480\pm 0.007$ $A_{s}(\mu K^{2})$ 2.400 $2.378\pm 0.175$ $2.378\pm 0.167$ $2.385\pm 0.167$ $\alpha_{s}$ -0.800 $0.777\pm 0.157$ $0.778\pm 0.148$ $0.771\pm 0.148$ $\beta_{s}$ -3.100 $3.102\pm 0.035$ $3.102\pm 0.028$ $3.101\pm 0.028$ $\epsilon_{ds}(\times 10^{2})$ 0 $015\pm 3.823 0.0$$11\pm 3.694 0.0$$010\pm 3.684$